OpenStudy (anonymous):

Okay, I need some help checking my answer to a function problem - Generate a function for the following: A number is equivalent to 515.85 and increases by 62 in one interval. How many intervals has passed for the number to reach 1795. The domain is any real number. – My answer is in the comment.

OpenStudy (anonymous):

\[x = 515.85 + 62(y)\]\[f(y) = 515.85 + 62y\]\[f(1) = 515.85 + 62(1)\]\[f(1) = 577.85\]\[.....\]\[f(20) = 1755.85\]

OpenStudy (anonymous):

I'm not done yet, still haven't reached 1795, just wondering if I got the rest correct...

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

please, solve this equation for y, then: \[1795=515.85+62*y\]

OpenStudy (anonymous):

20.62 intervals.

OpenStudy (michele_laino):

yes! you can round it to 21

OpenStudy (anonymous):

Alright, but how I presented the function is correct?

OpenStudy (michele_laino):

you can present your function as below: \[f(x)=515.85+62*x\]

OpenStudy (michele_laino):

so your guess it is correct!

OpenStudy (anonymous):

Thanks :D

OpenStudy (michele_laino):

Thanks! :)